Suppose the problem in Sample Problem 16.2 is projected at 24.5 m/s, 60° above the horizontal. Find the (a) time of flight, (b) range, and (c) maximum height.
1
Expert's answer
2020-08-25T11:13:08-0400

Given:

"v_0=24.5\:\rm m/s"

"\theta=60^{\circ}"

"g=9.8\:\rm m/s^2"

(a) the time of flight

"t=\frac{2v_0\sin\theta}{g}=\frac{2\times 24.5\sin 60^{\circ}}{9.8}=4.33\:\rm s"

(b) the range

"R=\frac{v_0^2\sin2\theta}{g}=\frac{24.5^2\sin 120^{\circ}}{9.8}=53\:\rm m"

(c) the maximum height

"H=\frac{v_0^2\sin^2\theta}{2g}=\frac{24.5^2\sin^260^{\circ}}{2\times 9.8}=23\:\rm m"